OpenStudy (anonymous):

Please Help if you can! You and your coworker together make $ 16 per hour. You know your coworker earns 10 percent more than you do. Your hourly wage is $7.62 . After taking Math your hourly wage is raised to $ 12. This is a raise of 57.49 %. After returning to work you can't help mentioning casually to your coworker that now you make__???????_ % more than he does.

OpenStudy (anonymous):

you + she = 16 she = you + 0.1*you (you + 01.you) + x*(you + 01.you) = 12 x = (12 - (you + 01.you))/(you + 01.you) * 100 (to get %)

OpenStudy (anonymous):

(12 - (7.62 + 0.762))/ (7.62 + 0.762) = 0.4316 = 43.16%

OpenStudy (anonymous):

ok thanks so much!!

OpenStudy (anonymous):

wait you = 7.62 you + she = 16 => she = 16 - 7.62 = 8.38 new you = 12 you now make = 12 - 8.38 = 3.62 you make 3.62/12 = 0.30166...*100 = 30.16% I must have some error!

OpenStudy (anonymous):

you + she = 16 she = you + 0.1*you she = 16 - you or she = you + 0.1*you x= (12 - she)/12 or x = (12 - you + 0.1*you)/12 *100% to get % both should give the same result. It is confusing that you have that you and the coworker make 16 and that the coworker makes 10% more. Why do you need both? This is solvable if you just know one or the other.

OpenStudy (anonymous):

or x = (12 - you - 0.1*you)/12 is the correct second option

OpenStudy (anonymous):

Final correct options x= (16 - you)/12 or x = (12 - you + 0.1*you)/12 I am tired. Too many mistakes...

OpenStudy (anonymous):

Final Final correct options (I am for real :) ) x= (16 - you)/12 or x = (12 - you - 0.1*you)/12 I am tired. Too many mistakes...

OpenStudy (anonymous):

The problem gives more information than is needed. She makes 8.38 as you said, so 12/8.38 = 1.43198, so he makes 43.20 percent more